Allison:

What is the approximate volume of a cylinder with a height of 3 ft and a radius of 4 ft? Use 3.14 to approximate pi. Express your answer as a decimal to the hundredths place. Enter your answer in the box. ft³

Allison:

@Hero

Allison:

@Elsa213 @SkyVoltage43

Allison:

@LoveJuliet112 @Falconmaster

LoveJuliet112:

ok the formula for the volume of a cylinger is 1.3 pi r 2 I think let me check

LoveJuliet112:

4/3rd pi r squared so it would be 4/3times pi times radius cubed

LoveJuliet112:

so ur answer is 267.9

SkyVoltage43:

V=πr2h=π·42·3≈150.79645

SkyVoltage43:

thats what i got

LoveJuliet112:

omg for a cylinder!!!

LoveJuliet112:

I did a sphere

SkyVoltage43:

lol so im right

Elsa213:

Led du questioner tew du wrong answer. Dun dunn dunnnn

LoveJuliet112:

volume for a cylinder is base time height

LoveJuliet112:

I believe the answer is 12

Elsa213:

3*4= 12

Elsa213:

I helped <3

LoveJuliet112:

No sky is right I just did the math

LoveJuliet112:

V=πr2h=π·42·3≈150.79645

Elsa213:

I did du multiplication.

Allison:

Okay never mind this question e.e

LoveJuliet112:

lol

LoveJuliet112:

He is right

Allison:

226.08...